OpenStudy (anonymous):

What is the value of x in the solution of this system? -3x-y+2z=7 2y+z=-5 3x+6y=-21 A) -4 B) -2 C) 1 D) 3

OpenStudy (anonymous):

have you tried ? or no clue ?

OpenStudy (anonymous):

Im not really sure how to do this? Can you explain it to me?

OpenStudy (anonymous):

you can use substitution or elimination

OpenStudy (anonymous):

for example .. if we go for substitution .. we can take the first equation and find what "3x" is : -3x-y+2z=7 then we get 3x=2z-y-7 we do so since in the third equation we have 3x.. so we may replace it with what we found: 3x+6y=-21 plug in here 3x and get: 2z-y-7+6y=-21 then you find here some information about z or y and plug it into the second equation..

OpenStudy (anonymous):

Im still a little confused what is the answer? maybe i can work my way backwards after i have it

OpenStudy (anonymous):

I would start by finding what z is equal to...using the second equation you can easily come to that answer. After that, take what you found for z and plug that into equation one and either solve for x or y (doesn't matter). after you have found that answer(either x or y) plug that into equation 3. Does that make sense?

OpenStudy (anonymous):

ah yes i understand now thanks everyone